OpenStudy (anonymous):

Integral of 1 / ( 1 + e ^ x) actual process please... :D

OpenStudy (anonymous):

Dang you got 103 fans now!

OpenStudy (anonymous):

Hi arman, Try looking at it like this\[\int\limits_{}{}\frac{1}{1+e^x}dx=\int\limits_{}{}\frac{1+e^x-e^x}{1+e^x}dx=\int\limits_{}{}1-\frac{e^x}{1+e^x}dx\]

OpenStudy (anonymous):

Haha, I know!

OpenStudy (anonymous):

\[=x-\log (1+e^x) + c\]

OpenStudy (anonymous):

log == ln always unless a base is made explicit

OpenStudy (anonymous):

It's an arctrig function and x-ln(1+e^x) is e^x/(1+e^x) (chain rule)

OpenStudy (anonymous):

II know that it's arctan of something ><

OpenStudy (anonymous):

The second integral makes use of this fact:\[\int\limits_{}{}\frac{f'(x)}{f(x)}dx=\log f(x)+c\]

OpenStudy (anonymous):

If f(x)=1+e^x, then f'(x)=e^x, which is the form you have after rearranging.

OpenStudy (anonymous):

I'm on my fone so I didn't see half of those equations till after I posted my last post.... lemmy try that... yummy and sec

OpenStudy (anonymous):

Gimmy a sec**

OpenStudy (anonymous):

u substitution then partial fraction decomposition

OpenStudy (anonymous):

I don't get why I can't think of these on my own -.-

OpenStudy (anonymous):

i got an answer of ln(e^x/(e^x+1))+c

OpenStudy (anonymous):

There's one way to check your answer - take the derivative and see if you get back to the integrand.

OpenStudy (anonymous):

yea it does

OpenStudy (anonymous):

just checked it

OpenStudy (anonymous):

Nope, pfd is in the chapter after this (I slipped it previousy and went ahead) so Lucas answer makes perfect sense...

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Good work

OpenStudy (anonymous):

Oh boy, integral of e^sqrtx from 0 to 4...

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

I don't even know what to use on this one...

OpenStudy (anonymous):

2(e^(sqrt(x))(sqrt(x)-1)|(0,4)

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

have u learned integration by parts yet?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

and i take it you know u substitution?

OpenStudy (anonymous):

start with your integral and set u equal to sqrt(x)

OpenStudy (anonymous):

Ok...

OpenStudy (anonymous):

u should get int(2u(e^u)du)

OpenStudy (anonymous):

do you follow me so far?

OpenStudy (anonymous):

You can put the dx in terms of u aswell? (Like I can write 2u instead of 2sqrtx? )

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

from this point you can use integration by parts

OpenStudy (anonymous):

Ok then that makes sense and i can use the uv vdu then.... -.- I'm retarded

OpenStudy (anonymous):

lol its hard to see that without substituting first tho

OpenStudy (anonymous):

No... I'm just retarded. And the answer is 2e^2

OpenStudy (anonymous):

did u evaluate at zero also?

OpenStudy (anonymous):

im getting 2(e^2)+2

OpenStudy (anonymous):

Good point, further testing my retardation.

OpenStudy (anonymous):

lol common mistake

OpenStudy (anonymous):

When you uv vdu sub, can you choose which term is your u and v? I'm pretty sure you can

OpenStudy (anonymous):

usually it depends on what you have

OpenStudy (anonymous):

to remember what to take for u i use "lipet"

OpenStudy (anonymous):

which gives you an order in which to take your u (Lograthims Inverse.trg Polynomials Exponetials Trigonometrics)

OpenStudy (anonymous):

Well I got a badass equation here: integral of (lnx) /x^2 from 1 to 2

OpenStudy (anonymous):

I know I gotta use uv vdu on it

OpenStudy (anonymous):

judging by lipet take log to be your u

OpenStudy (anonymous):

and 1/x^2 to be your dv

OpenStudy (anonymous):

im getting (-ln(x)/x)-(1/x)+c

OpenStudy (anonymous):

your evaluated integral should come out to be (-ln(2)/2)+1/2

OpenStudy (anonymous):

were you able to solve it?